two point charges are located on the positive x axis of a coordinate system charge q1=1.0nC is 2.0cm from the origin, another charge q2=-3.0nC os 4.0cm from the origin .what is the total force exerted by these two charges on charge q3=5.0nC located at The origin

1
Expert's answer
2022-03-16T18:30:12-0400

Force

"F_1=\frac{kq_1q_2}{r^2}"

"F=\frac{kq_1q_3}{r^2}+\frac{kq_2q_3}{r^2}"

"F={kq_3}{r^2}"

"F=kq_3(\frac{q_1}{r^2}+\frac{q_2}{r^2})"

"F=9\times10^9\times5\times10^{-9}(\frac{1\times10^{-9}}{0.02^2}-\frac{3\times10^{-9}}{0.04^2})"

"F=9\times10^9\times5\times10^{-18}(\frac{1}{0.02^2}-\frac{3}{0.04^2})"

"F=9\times10^9\times5\times10^{-18}\times10^4(\frac{1}{2^2}-\frac{3}{4^2})"

"F=4.5\times10^{-4}\times\frac{4}{16}=1.125\times10^{-4}N"